40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Mar 5 06:27:14 2013, yuta, Update, LSC, intra-cavity power dependence on mirror misalignment PRM_PIT_20130305a.pngPRM_YAW_20130305b.pngITMX_YAW_20130305.pngITMY_YAW_20130305.png
    Reply  Tue Mar 5 23:00:08 2013, yuta, Update, LSC, Yarm and PRC g-factor from misalignment measurement ITMY_YAW_20130305_DC.pngPRM_PIT_20130305a_DC.pngPRM_YAW_20130305b_DC.png
       Reply  Tue Mar 5 23:37:11 2013, yuta, Update, SUS, PRM angular motion spectra PRMangularmotion.png
    Reply  Wed Mar 6 02:46:20 2013, Manasa, Update, LSC, PRMI locking for g-factor measurement POPDCspec_PRMIlocked.png
       Reply  Wed Mar 6 09:44:29 2013, Koji, Update, LSC, PRMI locking for g-factor measurement 
       Reply  Fri Mar 8 21:00:05 2013, Manasa, Update, PSL, PMC fixed 
Message ID: 8263     Entry time: Fri Mar 8 21:00:05 2013     In reply to: 8237
Author: Manasa 
Type: Update 
Category: PSL 
Subject: PMC fixed 

Quote:

  Mystery

1. Filter module (FM1) on PRCL and MICH show significant delay while enabling and disabling.

2. I tried to fix PMC alignment (PMC trans was 0.76). I was not able to get PMC trans more than 0.79.
PMC has been this way since yesterday.

3. MICH is still bright when locked (ASDC_OUT reads 0.08 for dark and 2.0 for bright). We suspect it is because of the AS55_I error offset that persists even after running LSCoffsets script.

4. PRMI shows some dither at 3Hz when locked.

 [Koji, Manasa]

PMC is fixed with 0.84 in transmission. It was misaligned in pitch (fixing this increased PMC_trans to 0.822 from 0.773) and Koji also touched the wave plate and polarizer (changed PMC_trans to 0.845).

ELOG V3.1.3-